if 8x - 4 = 6x - 10 find the value of 5x show work please!
1
Expert's answer
2012-11-27T07:47:47-0500
8x - 4 = 6x - 10
8x-6x=4-10
2x=-6
x=-3
Now let's find required value: 5x=5*(-3)=-15